# PointSAM: Pointly-Supervised Segment Anything Model for Remote Sensing Images

### v1

@@ IV-E1 Visualization of the NPC Process
To visually demonstrate the effect of NPC during training, we present the results from the initial mask to the refined mask across three datasets. As shown in Fig. 8, the red points and green points represent positive prompts and negative prompts, respectively. In the first row for the NWPU VHR-10 dataset, the texture information of the tennis court is quite subtle, causing the initial mask to include an adjacent tennis court. After incorporating NPC, overlapping objects are treated as negative prompts, leading to the removal of excess masks. As the number of positive prompts increases, prompts located at the edges of objects are more likely to cause semantic ambiguity. This ambiguity can be accurately eliminated through NPC. For the buildings (second row) in the WHU dataset, they all have similar colors, which makes the mask indicated by the point prone to interference. Thanks to the large number of buildings in each image, NPC can easily locate nearby ambiguous masks, thus constraining the target mask. For the most challenging HRSID-inshore dataset, due to the SAR imaging mechanism, the color of each ship and the inshore scene appear identical (third row). Moreover, the targets are small and may be hollow. Therefore, if each negative prompt is constrained, a large amount of non-target regions will be designated as the mask. It can be seen that our method effectively suppresses redundant regions, regardless of the number of prompts.
@@ IV-E2 Visualization of results from different methods
We then present the comparative results of different methods across various datasets. In the Fig. 9, rows 1-3 show the results on NWPU VHR-10 with 1 to 3 prompts; similarly, rows 4-6 display the results on WHU, and rows 7-9 show the results on HRSID-inshore. It can be observed that due to the bird’s-eye view in remote sensing images, there is a significant difference from natural images. Directly using the original SAM leads to an inability to distinguish each target clearly. For example, in the sixth row, the white building on the left and the parking lot on the right are treated as the same object. Even more notably, in the HRSID-inshore dataset, most of the inshore area is wrongly labeled as the target mask. Self-training transfers the source model to the target data, reducing more redundant areas and producing relatively more complete predicted masks compared to direct testing. However, it still fails to mitigate the interference between adjacent objects, such as the tennis court in the second row and the building in the fifth row. DePT, Tribe, and WeSAM are all improvements based on self-training, and they handle mask details better than self-training. However, in more challenging scenarios, they still fail to achieve optimal results. For example, in the third row, the storage tank and its shadow are not separated, and the ships in the inshore scene are not accurately segmented (rows 7-9). In contrast, our method excels at handling objects in dense scenes, achieving performance close to the ground truth.
@@ IV-F PointSAM as a Detection Box Generator
In this section, we serve PointSAM as a point-to-box generator. PointSAM can generate corresponding masks based on points, and by calculating the minimum enclosing rectangle of the mask, we can obtain the corresponding horizontal bounding box (HBB). These HBBs can then be fed into a detector that converts horizontal boxes to rotated boxes, achieving point-supervised oriented object detection. To validate the effectiveness of this approach, we conducted experiments on the HRSID dataset, which includes both inshore and offshore scenarios. All experiments were conducted with an input size of 800×800, running for 12 epochs, and using ResNet-50 as the backbone. As shown in Table 6, we compared our method with representative algorithms based on OBB supervision, HBB supervision, and point supervision. It can be observed that the H2RBox-v2 and the method proposed by Yue et al. [71] based on HBB can achieve performance comparable to OBB supervision. The poor performance of H2RBox may be attributed to the large number of small objects in the HRSID dataset. Therefore, our approach also utilizes H2RBox as the detector for converting HBB to OBB. Compared to vanilla SAM, our method achieves a 15% improvement. This is because directly using SAM can result in unclear segmentation masks for objects in dense scenes, which in turn leads to inaccuracies in the minimum enclosing rectangles. Similarly, our method slightly outperforms Point2Rbox. Essentially, both Ours and Point2RBox leverage prior knowledge to learn the size information of the targets. There remains a gap of nearly 20% compared to the HBB-supervised methods. Future work could focus on integrating multiple types of priors to bridge this gap.
TABLE VI: Comparisons results of different detectors based on HRSID.

### v2

@@ IV-E1 Visualization of the NPC Process
To visually demonstrate the impact of NPC during training, we present the progression from initial masks to refined masks across three datasets, as shown in Fig. 8. Red and green points denote positive and negative prompts, respectively. In the NWPU VHR-10 dataset, the subtle texture of tennis courts causes initial masks to cover adjacent courts. Positive prompts near object edges, such as cars, are often affected by nearby objects. By treating overlapping targets as negative prompts, redundant masks are effectively removed. As the number of positive prompts increases, prompts near object boundaries, such as airplanes or harbors, tend to cause semantic ambiguity. This ambiguity is resolved by NPC. In the WHU dataset, buildings with similar colors often result in interference in masks generated from prompts. Given the high density of buildings in each image, the method identifies ambiguous masks nearby and refines the boundaries of the target mask. The HRSID inshore dataset is the most challenging due to the SAR imaging mechanism. Ships and the surrounding scenes share nearly identical colors. Additionally, the targets are small and may have hollow structures. If constraints are applied to each negative prompt, non-target regions are likely to be included in the mask. Despite these challenges, the method suppresses redundant regions effectively, regardless of the number of prompts.
@@ IV-E2 Visualization of results from different methods
We then present the comparative results of different methods across various datasets in Fig. 9, Fig. 11, and 10. Rows 1-3 show the results with 1 to 3 prompts. It can be observed that due to the bird’s-eye view in remote sensing images, there is a significant difference from natural images. Directly using the original SAM leads to an inability to distinguish each target clearly. For example, in the third row in Fig.10, the white building on the left and the parking lot on the right are treated as the same object. Even more notably, in the HRSID-inshore dataset (see Fig.11), most of the inshore areas are misidentified as positive masks. Self-training transfers the source model to the target data, reducing more redundant areas and producing relatively more complete predicted masks compared to direct testing. However, it still fails to mitigate the interference between adjacent objects, such as the tennis court in the second row in Fig. 9 and the building in the fifth row in Fig. 10. DePT, Tribe, and WeSAM are all improvements based on self-training, and they handle mask details better than self-training. However, in more challenging scenarios, they still fail to achieve optimal results. For instance, in the third row of Fig. 9, the storage tank and its shadow remain indistinguishable, and ships in inshore scenes are not accurately segmented (Fig. 11). In contrast, our method excels at handling objects in dense scenes, achieving performance close to the ground truth.
@@ IV-F PointSAM as a Detection Box Generator
In this section, we serve PointSAM as a point-to-box generator. PointSAM can generate corresponding masks based on points, and by calculating the minimum enclosing rectangle of the mask, we can obtain the corresponding horizontal bounding box (HBB). These HBBs can then be fed into a detector that converts horizontal boxes to rotated boxes, achieving point-supervised oriented object detection. To validate the effectiveness of this approach, we conducted experiments on the HRSID dataset, which includes both inshore and offshore scenarios. All experiments were conducted with an input size of 800×800, running for 12 epochs, and using ResNet-50 as the backbone. As shown in Table 6, we compared our method with representative algorithms based on OBB supervision, HBB supervision, and point supervision. It can be observed that the H2RBox-v2 and the method proposed by Yue et al. [81] based on HBB can achieve performance comparable to OBB supervision. The poor performance of H2RBox may be attributed to the large number of small objects in the HRSID dataset. Therefore, our approach also utilizes H2RBox-v2 as the detector for converting HBB to OBB. Compared to vanilla SAM, our method achieves a 15% improvement. This is because directly using SAM can result in unclear segmentation masks for objects in dense scenes, which in turn leads to inaccuracies in the minimum enclosing rectangles. Similarly, our method slightly outperforms Point2Rbox. Essentially, both Ours and Point2RBox leverage prior knowledge to learn the size information of the targets. There remains a gap of nearly 20% compared to the HBB-supervised methods. Future work could focus on integrating multiple types of priors to bridge this gap.
TABLE VII: Comparisons results of different detectors based on HRSID.